OpenStudy (anonymous):

solve the system of equations using the matrix tool.(explanation please) 4x+5y=13 8x-2y=14

OpenStudy (anonymous):

\[write \in the form AX=B\] \[Find \left| A \right|,then adjoint A\] \[inverseA=\frac{ adj A }{ \left| A \right|}\] X=inverse of A* B

OpenStudy (anonymous):

\[A=\left[\begin{matrix}4 &5 \\ 8 & -2\end{matrix}\right]\] \[X=\left(\begin{matrix}x \\ y\end{matrix}\right)\] \[B=\left(\begin{matrix}13 \\ 14\end{matrix}\right)\]\[\left| A \right|=\left| \begin{matrix}4 & 5 \\8 & -2\end{matrix} \right|=4\left( -2 \right) -5\left( 8 \right) =-8-40=-48\neq 0 \] hence inverse of A exists.

OpenStudy (anonymous):

Find matrix of of co factors of each element of each row then adj. of A is the transpose of above matrix. i think now you can solve it.

OpenStudy (anonymous):

that was so confusing

OpenStudy (anonymous):

cofactors of elements of first row of \left| A \right|are -2,-8 cofactors of elements of second row of \left| A \right| are -5,4 \[adj. A=\left[\begin{matrix}-2 & -8 \\ -5 & 4\end{matrix}\right]\prime \] \[adj.A=\left[\begin{matrix}-2 & -5 \\ -8 & 4\end{matrix}\right]\] \[A ^{-1}=\frac{ 1 }{ -48 }\left[\begin{matrix}-2 &-5 \\ -8 & 4\end{matrix}\right]\] \[X=A ^{-1}B\] Now try to find X

OpenStudy (anonymous):

Let me know if you can solve.

OpenStudy (anonymous):

cofactors of elements of first row of | A| are-2,-8

OpenStudy (anonymous):

\[X=\frac{ -1 }{48 }\left[\begin{matrix}-2 & -5\\ -8 & 4\end{matrix}\right]\left(\begin{matrix}13 \\14\end{matrix}\right)\] \[X=\frac{ -1 }{ 48 }\left(\begin{matrix}-2*13+(-5)*14 \\ -8*13+4*14\end{matrix}\right)\] \[X=\frac{ -1 }{ 48 }\left(\begin{matrix}-26-70 \\ -104+56\end{matrix}\right)\] \[X=\frac{ -1 }{48 }\left(\begin{matrix}-96 \\-48\end{matrix}\right)=\left(\begin{matrix}-96/-48 \\ -48/-48\end{matrix}\right)\] \[\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}2 \\ 1\end{matrix}\right)\] x=2,y=1

OpenStudy (anonymous):

\[\left| A \right|=\left| \begin{matrix}a & b \\ c &d\end{matrix} \right|\] \[co factor of a=\left( -1 \right)^{ri+j}d=\left( -1 \right)^{1+1}d=\left( -1 \right)^{2}d=d\] i=no. of row j=no. of column \[similarly cofactor of b=\left( -1 \right)^{i+j}c=\left( -1 \right)^{i+j}c=\left( -1 \right)^{1+2 }c=-c\] \[cofactor of c=\left( -1 \right)^{i+j}b=\left( -1 \right)^{2+1}c\] \[and cofactor of d=\left( -1 \right)^{i+j}a=\left( -1 \right)^{2+2}a=a\]